What force must be exerted on the master cylinder of a hydraulic lift to support the weight of a 2 500 kg car a large car resting on the slave cylinder? The master cylinder has a 4.00 cm diameter and the slave has a 25.0 cm diameter.

1
Expert's answer
2022-01-27T13:17:59-0500

Gives

"m_2=2500kg"

"r_1=0.02m"

"r_2=\frac{25\times10^{-2}}{2}" =0.125m

"P_1=P_2"

"\frac{F_1}{A_1}=\frac{F_2}{A_2}"

"F_1=\frac{A_1}{A_2}\times F_2"

"F_1=\frac{r_1^2}{r_2^2}\times F_2"


"F_1=\frac{0.02^2}{0.125^2}\times2500\times9.8=627.2N"